6e^(1-x)+1=25, find x.
Answer by sdmmadam@yahoo.com(530)

6e^(1-x)+1=25, find x.
6e^(1-x)+1=25----(1)
6e^(1-x) = 25-1
6e^(1-x) = 24
e^(1-x)=24/6
e^(1-x) = 4
log[e^(1-x)=loge(4)
(1-x)loge(e) = loge(4)
(1-x)X1 = loge(4)
1-x= loge(4)
1-loge(4) = x
Answer: x = [1-loge(4)]
Verification: Putting x = [1-loge(4)] in (1),
LHS = 6e^(1-x)+1
=6e^[1-(1-log4)] +1 (here the base is e)
=6e^(1-1+log4)+1
=6e^(log4)+1
=6X4 + 1 (since e^(log(p)) = p )
=24+1 = 25 = RHS
Therefore our value for x is correct